OpenStudy (anonymous):

Who knows how to convert radians to degrees ? For values greater than 360, find the equivalent angle less than 360°.

OpenStudy (mathmale):

The "radians to degrees" conversion factor is \[\frac{ 180~\deg }{ \pi~rad }\]

OpenStudy (anonymous):

But find the equivalent value , how do I do that ?

OpenStudy (mathmale):

Example: Convert 2.1 rad to degrees:\[\frac{ 2.1~rad }{ 1 }\frac{ 180~\deg }{ \pi~rad }\]

OpenStudy (mathmale):

Reduce as far as you can. Use pi = 3.14.

OpenStudy (anonymous):

I don't understand this , look it says convert 11*pi over 4 radians to degrees . For values greater than 360 , find the equivalent angle less than 360° .

OpenStudy (anonymous):

pi rad = 180 degree 1 rad = 180/pi degree for the values greater than 360 degree,say 3pi rad = 540 degree u should subtract 540 from 360,and u'll get the equivalent angle

OpenStudy (anonymous):

By equivalent angle, they mean \(361^\circ\) is equivalent \(1^\circ\), and so on.

OpenStudy (anonymous):

Huh ?^

OpenStudy (anonymous):

Do you know about the unit circle?

OpenStudy (anonymous):

I don't know anything about this , I just started the lesson today , im so confused .

OpenStudy (solomonzelman):